OpenStudy (anonymous):

Please help! Will medal and fan! Kelly needs to order lunch for orders 6 people at a business meeting. Her menu choices are chicken salad for a cost of $5 per person and egg salad for a cost of $4 per person. She only has $28 to spend. More people want chicken salad 1. Write and graph one equation in a system for this situation. 2. Write a second equation in a system. 3. What do x and y represent? 4. How many of each type of lunch can she order?

OpenStudy (anonymous):

3. x= 82 y=19

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

x= 82 y=19 Simple.

OpenStudy (anonymous):

Ok. So say Kelly orders 1 chicken salad, then it means she ordered 5 egg salads. If she ordered \(x\) chicken salads, then she'll have to order the rest, which is \((6-x)\) egg salads. For first equation, let's set \(y\) to be the meal's price. It consist of the price of all chicken salads ordered and the price of all egg salads ordered: $$ y = 5x + 4(6-x) $$ Again, \(x\) represents number of chicken salads ordered, and \(y\) represents the meal's price. A second equation would come to limit the price to the money Kelly has: $$ y \le 28 $$ Notice there are more constraints to add on the values (\(x\) cannot be negative, no such thing negative number of salads or fractional number of salads) so just keep that in mind. Now, since we know the price based on the number of chicken salads, we can say for what number of salads it matches the second restriction: $$ 5x + 4(6-x) \le 28 \\ 5x +24 -4x \le 28 \\ x \le 28 - 24 \\ x \le 4 $$ So she can order at max 4 chicken salads (which will come with 2 egg salads).

OpenStudy (anonymous):

@pitamar great answer/description! ;) Medal!!!!

OpenStudy (anonymous):

Thanks =)

OpenStudy (anonymous):

Sure thing @pitamar

OpenStudy (anonymous):

Thanks for that.